# LEMON-Mapping: Loop-Enhanced Large-Scale Multi-Session Point Cloud Merging and Optimization for Globally Consistent Mapping

### v2

@@ Abstract
Multi-robot collaboration is becoming increasingly critical and presents significant challenges in modern robotics, especially for building a globally consistent, accurate map. Traditional multi-robot pose graph optimization (PGO) methods ensure basic global consistency but ignore the geometric structure of the map, and only use loop closures as constraints between pose nodes, leading to divergence and blurring in overlapping regions. To address this issue, we propose LEMON-Mapping, a loop-enhanced framework for large-scale, multi-session point cloud fusion and optimization. We re-examine the role of loops for multi-robot mapping and introduce three key innovations. First, we develop a robust loop processing mechanism that rejects outliers and a loop recall strategy to recover mistakenly removed but valid loops. Second, we introduce spatial bundle adjustment for multi-robot maps, reducing divergence and eliminating blurring in overlaps. Third, we design a PGO-based approach that leverages refined bundle adjustment constraints to propagate local accuracy to the entire map. We validate LEMON-Mapping on several public datasets and a self-collected dataset. The experimental results show superior mapping accuracy and global consistency of our framework compared to traditional merging methods. Scalability experiments also demonstrate its strong capability to handle scenarios involving numerous robots.
Index Terms:
Multi-robot mapping, map merging, SLAM.I Introduction
@@
In addition to the above mapping approaches, to achieve larger-scale map reconstruction and multi-robot collaboration, several multi-robot SLAM systems [27, 28] and offline map fusion techniques [29, 30, 31, 32, 33] have shown promising results in recent years. In general, these methods focus on two aspects: removing outliers in loop closure data and achieving map fusion through multi-robot PGO. However, PGO-based methods are fundamentally the same in single-robot and multi-robot applications as they both use odometry constraints and loop closure constraints for map refinement. As discussed in the previous paragraph, these approaches neither directly utilize the geometric information in multi-robot maps nor fully leverage loop closure data. Consequently, they only ensure the basic alignment and consistency of the global map but fail to generate high-quality point cloud maps. In particular, they often suffer from misalignment in regions with overlap between robots, as demonstrated in Fig. 8 (b) and (c).
@@ II-B Multi-Session Map Merging
Multi-session map merging aims to integrate submaps from multiple agents either with or without initial pose estimates into a unified, globally consistent map. SegMap [38] extracts semantic features from 3D point clouds to estimate 6-DoF transformations and utilizes incremental pose graph optimization to achieve map fusion. However, it heavily depends on accurate semantic segmentation, limiting its applicability to structured urban environments. AutoMerge [30] introduces a city-scale merging framework that avoids relying on initial alignment, but its performance degrades in complex environments due to the lack of elimination of loop outliers. PCM [31, 32, 33] selects reliable loop sequences through pairwise consistency maximization, but struggles on datasets with high loop error rates. It is also severely affected by the accumulated drift of odometry in large-scale environments. LAMM [29] enhances place recognition by removing dynamic objects through M-Detector [39] and using the robust loop detection method BTC [20], yet it only merges map by pose graph optimization, which may suffer from local divergence in the overlapping areas of the multi-robot map.
@@ V-A DBA for Isolated Loops
Our DBA is built upon BALM2 [24]. However, it relies on accurate initial plane estimation, which may fail to converge reliably when the same plane observed by different robots appears misaligned due to map divergence. In such cases, the same structure may be voxelized as separate multi-layer planes (shown in Fig. 3 (a)), making direct optimization invalid. To address this, we partition the poses within the spatial window by robot into distinct clusters, each with its own point cloud, as illustrated in the left part of Fig.3. We then apply GICP [42] to roughly align these clusters, reducing plane stratification and enabling reliable plane estimation for subsequent optimization. After the initial pose graph optimization incorporating an isolated loop, the poses near the loop center are effectively aligned and optimized. However, as the distance from the loop center increases, the influence of loop constraints in FPGO gradually weakens, leading to less accurate optimization and alignment of distant poses. Therefore, for an isolated loop, the key point is to figure out how large its influence range is and align the surrounding affected poses. We develop DBA to tackle this issue.
@@
For simplicity, we use the following notation, and we refer the reader to [24] for more detailed information. In this paper, feature is seen as point clusters, denoted by set , and the corresponding point cluster coordinate is defined as:
A BA formulation with determining LiDAR poses , where is the number of poses, and feature parameters , where is the number of features, starts with the optimization problem:
@@
In DBA, we categorize the poses into groups based on their distances from the loop. Each group contains a set of poses. The number of poses in the -th group is denoted as (for ). In the following part, we first demonstrate that in this incremental form of BA, the optimal update formulation remains the same as performing BA jointly on all features. We then prove that the algorithmic complexity is slightly smaller compared to the traditional BA methods, and our approach achieves a relatively higher confidence level which should perform better in practice.
@@
Lemma 1.
If is a continuous real symmetric matrix function (i.e., is continuous in and is symmetric for every ), then its eigenvalue function is also continuous.
@@
See Appendix (Appendix: PROOF OF LEMMA 1). ∎
According to the definition of matrix , we could obviously get . Hence, based on Lemma 1, we could easily get every item of cost function is a continuous function. In the formulation of DBA, cost item could be rewritten as:
where constant . While the cost function is continuous, according to the chain rule, the gradient would be:
where it does not change when optimizing compared with jointly solving the BA. The same is true for the second-order gradient. Based on (5), we only need to freeze the update for poses ().
The solver that jointly solves this BA problem has a complexity of , while by masking the Jacobian and Hessian flexibly, our DBA could achieve a complexity of:
@@
Assumption 1.
Small range LiDAR poses that participate in the BA have smaller covariance in measurement noise compared to LiDAR poses in wider ranges.
According to Assumption 1, with the previous group categorization, define and , we can obtain that:
Thus, under our assumption, fixing the number of iterations yields lower measurement noise, while fixing the confidence level leads to faster convergence.
@@ VI-A First Pose Graph Optimization
Following loop processing, we perform the first pose graph optimization (FPGO) to coarsely align the trajectories of all robots. We construct a centralized pose graph shown in Fig. 5, in which each robot’s pose sequence is included as a subgraph. The graph incorporates three types of constraint: odometry, self-loops, and inter-loops. The first pose of the first robot is fixed to define the origin point of the world coordinate, ensuring that all other robot trajectories are aligned to this global reference frame. With this initial alignment established, we acquire a rough estimate of the relative positions among all robot trajectories, which enables us to recover the valid but previously rejected loops (in Section IV).
@@ VII-A Experimental Setup
We first evaluate our proposed spatial bundle adjustment by comparing it with the conventional LiDAR-based bundle adjustment methods BALM2 [24] and HBA [25] (Section VII-B). Next, we assess the performance of our complete multi-session merging framework against DCL-SLAM [27] and LAMM [29] (Section VII-C). An ablation study is then conducted to analyze the contributions of individual components within our framework (Section VII-D). Finally, we test the scalability of our method in handling a large number of sessions (Section VII-E).
@@ VII-B Single-Robot Study
VII-C Multi-Robot Study
@@
VII-D Ablation Study
@@
Table V shows the number of loops in different loop processing stages, as well as the results of the first pose graph optimization with and without loop recall. The experimental results show that the first PGO achieves smaller errors and better performances when mistakenly rejected loops are recalled. Beyond its impact on PGO, loop recall also introduces additional constraints that benefit subsequent bundle adjustment.
@@
Table VI summarizes the results, with the best values highlighted in bold. The complete framework consistently achieves the lowest RMSE, validating the effectiveness of combining spatial BA and two PGO steps. For the variant of FPGO + BA, although local BA improves relative accuracy, it may disrupt the continuity of odometry since it only refines loop regions. Consequently, its standalone use can increase RMSE compared to FPGO alone. However, the last PGO integrates both local BA constraints and odometry continuity, propagating local accuracy to the global map and significantly reducing RMSE.
@@
VII-E Scalability Study
@@
Table VII summarizes the corresponding results of three cases. Our framework achieves a success rate 100% in all experiments, including the five-session, ten-session, and twenty-session scenarios. These results demonstrate the scalability and robustness of the LEMON-Mapping system when dealing with large numbers of multi-robot map merging.
@@ VIII Conclusion
Appendix: PROOF OF LEMMA 1
@@
A standard result on the continuous dependence of polynomial roots on coefficients now implies that each ordered root varies continuously with the coefficients , and hence with itself. Therefore, every eigenvalue function is continuous.

### v3

@@ Abstract
Multi-robot collaboration is critical but challenging in building globally consistent maps. Traditional multi-robot pose graph optimization (PGO) methods ensure basic global consistency but ignore the geometric structure of the map, and only use loop closures as constraints between pose nodes, leading to divergence and blurring in overlapping regions. To address this, we propose LEMON-Mapping, a loop-enhanced framework for large-scale, multi-session point cloud fusion and optimization. We re-examine the role of loops in multi-robot mapping and present three key innovations. First, we develop a robust loop processing mechanism with outlier rejection and a recall strategy to recover valid loops. Second, we introduce spatial bundle adjustment to reduce divergence and eliminate blurring in overlapping areas. Third, we design a PGO-based optimization that integrates refined bundle adjustment constraints to propagate local accuracy globally. Experiments on public and self-collected datasets demonstrate that LEMON-Mapping achieves superior accuracy, consistency, and scalability over traditional approaches in large-scale multi-robot scenarios.
Note to Practitioners
\@IEEEabskeysecsize\@IEEEgobbleleadPARNLSPIn this paper, we address inaccurate and inconsistent large-scale multi-robot map fusion, where conventional multi-robot SLAM ignores geometric structures, causing divergence and blurring in overlapping regions. In real-world exploration, multi-robot mapping is designed to extend coverage, but inconsistent and inaccurate map fusion degrades map quality and affects downstream applications like re-localization, navigation, and robotic operation, potentially leading to navigation failures and unsafe behaviors. LEMON-Mapping integrates robust loop processing, spatial bundle adjustment, and pose graph optimization to reconstruct accurate and globally consistent multi-robot maps, providing reliable support for multi-robot collaboration, localization, and autonomous navigation.
I Introduction
@@
In addition to the above mapping approaches, to achieve larger-scale map reconstruction and multi-robot collaboration, several multi-robot SLAM systems [27, 28] and offline map fusion techniques [29, 30, 31, 32, 33, 34] have shown promising results in recent years. In general, these methods focus on two aspects: removing outliers in loop closure data and achieving map fusion through multi-robot PGO. However, PGO-based methods are fundamentally the same in single-robot and multi-robot applications as they both use odometry constraints and loop closure constraints for map refinement. As discussed in the previous paragraph, these approaches neither directly utilize the geometric information in multi-robot maps nor fully leverage loop closure data. Consequently, they only ensure the basic alignment and consistency of the global map but fail to generate high-quality point cloud maps. In particular, they often suffer from misalignment in regions with overlap between robots, as demonstrated in Fig. 8 (b) and (c).
@@ II-B Multi-Session Map Merging
Multi-session map merging aims to integrate submaps from multiple agents either with or without initial pose estimates into a unified, globally consistent map. Some multi-robot exploration frameworks such as SMMR-Explore [39] and MR-GMMExplore [40] address multi-robot exploration and map fusion under communication limits. However, the former is limitted to 2D point cloud submaps while the latter heavily suffers from loss of geometric details caused by Gausssian Mixture Models (GMMS) based submaps. And their lack of pose graph optimization (PGO) prevents them from achieving global consistency. Therefore, these multi-robot exploration frameworks are both unsuitable for generation of large-scale geometric accurate point cloud map. SegMap [41] extracts semantic features from 3D point clouds to estimate 6-DoF transformations and utilizes incremental pose graph optimization to achieve map fusion. However, it heavily depends on accurate semantic segmentation, limiting its applicability to structured urban environments. AutoMerge [30] introduces a city-scale merging framework that avoids relying on initial alignment, but its performance degrades in complex environments due to the lack of elimination of loop outliers. PCM [31, 32, 33] selects reliable loop sequences through pairwise consistency maximization, but struggles on datasets with high loop error rates. It is also severely affected by the accumulated drift of odometry in large-scale environments. LAMM [29] enhances place recognition by removing dynamic objects through M-Detector [42] and using the robust loop detection method BTC [20], yet it only merges map by pose graph optimization, which may suffer from local divergence in the overlapping areas of the multi-robot map.
@@ V-A DBA for Isolated Loops
Our DBA is built upon BALM2 [24]. However, it relies on accurate initial plane estimation, which may fail to converge reliably when the same plane observed by different robots appears misaligned due to map divergence. In such cases, the same structure may be voxelized as separate multi-layer planes (shown in Fig. 3 (a)), making direct optimization invalid. To address this, we partition the poses within the spatial window by robot into distinct clusters, each with its own point cloud, as illustrated in the left part of Fig.3. We then apply GICP [45] to roughly align these clusters, reducing plane stratification and enabling reliable plane estimation for subsequent optimization. However, GICP only provides a coarse alignment of multi-robot submaps rather than true fine fusion. It simply adjusts the relative transformation of the assembled point cloud submaps of different pose sequence, but it does not optimize multi-robot poses at the frame level or refine point cloud at the single scan level. After the initial pose graph optimization incorporating an isolated loop, the poses near the loop center are effectively aligned and optimized. However, as the distance from the loop center increases, the influence of loop constraints in FPGO gradually weakens, leading to less accurate optimization and alignment of distant poses. Therefore, for an isolated loop, the key point is to figure out how large its influence range is and align the surrounding affected poses. We develop DBA to tackle this issue.
@@
For simplicity, we use the following notation, and we refer the reader to [24] for more detailed information. In this paper, feature is seen as point clusters, the point cluster for the -th feature is denoted by set , where is the number of poses, and the corresponding point cluster coordinate is defined as:
A BA formulation with determining LiDAR poses , and feature parameters , where is the number of features, starts with the optimization problem:
@@
In DBA, we categorize the poses into groups based on their distances from the loop. Each group contains a set of poses. The number of poses in the -th group is denoted as (for ). In the following part, we first demonstrate that in this incremental form of BA, the optimal update formulation remains the same as performing BA jointly on all features. We then prove that the algorithmic complexity is smaller compared to the traditional BA methods, and our approach achieves a relatively higher confidence level which should perform better in practice.
@@
Assumption 1.
Small range LiDAR poses that participate in the BA have significantly smaller measurement noise covariance than that associated with the outer pose .
Lemma 1 (Continuity in Eigenvalue Functions).
If is a continuous real symmetric matrix function (i.e., is continuous and is symmetric for every ), then its eigenvalue function is also continuous.
@@
See Appendix (VIII-A). ∎
According to the definition of matrix , we could obviously get . Hence, based on Lemma 1, we could easily partition conformably the LM Jacobian and Hessian for DBA as:
Using LM optimization in Eq. 5, we could respectively derive the update of pose for group for the joint BA and DBA:
where is the Schur’s complement.
The joint optimization can be viewed as an update method that refines the coupled terms for the DBA method. Hence from Eq. 7, we could define two refinement rates about the Hessian () and the Jacobian () as ratio of the coupling contribution to the main term:
The smaller these two refinement rates are (i.e., closer to ), the less significant the coupled terms become relative to the main terms. According to Assumption 1, consequently, the information matrix (Hessian approximation) , which represents the certainty of the inner poses , is assumed to be strong (i.e., possessing large singular values), formally implying that the norm of its pseudo-inverse is relatively small:
which in turn drives and towards zero. Consequently, the bias of the two method is minimized, and the high-speed DBA method closely approximates the high-accuracy joint optimization of traditional BA.
Lemma 2 (Computational Complexity Reduction of DBA).
Let the total number of poses be , where denotes the number of poses in the -th diffusion group, for simplicity. A joint BA optimization over all poses has complexity
In contrast, DBA performs BA incrementally and its total complexity is
which is strictly no larger and generally substantially smaller. In particular, when the diffusion groups are of comparable size, i.e., , the quadratic and cubic terms are reduced by factors of approximately and , respectively.
Proof.
See Appendix (VIII-B). ∎
The above proof uses a maximum of diffusion steps, while in practice, DBA is used when having a strong confidence of inner group poses for a small number of times far below .
@@
Lemma 3 (Covariance ordering: DBA vs joint BA).
Similar to the partition above, let the Jacobian-derivative blocks w.r.t. feature clusters be
Define the joint-pose covariance perturbation due to cluster noise as
and let be its bottom-right block (covariance of block 1 under joint BA). For DBA (block 0 frozen, optimizing only block 1) define
Then, under the usual invertibility assumptions on and , the following PSD ordering holds:
Consequently, freezing well-measured inner poses (block 0) yields a covariance for the active poses that is no larger (and typically smaller) than the covariance those same poses would have under a full joint BA. Note that the inequality follows without assuming inner poses are better measured.
Proof.
See Appendix (VIII-C). ∎
@@ VI-A First Pose Graph Optimization
Following loop processing, we perform the first pose graph optimization (FPGO) to coarsely align the trajectories of all robots. We construct a centralized pose graph shown in Fig. 5, in which each robot’s pose sequence is included as a subgraph. The graph incorporates three types of constraint: odometry, self-loops, and inter-loops. The first pose of the first robot is fixed to define the origin point of the world coordinate, ensuring that all other robot trajectories are aligned to this global reference frame. With this initial alignment established, we acquire a rough estimate of the relative positions among all robot trajectories, which enables us to recover the valid but previously rejected loops (in Section IV). The recalled loops are added back to the correct set of loops and used as constraints in the FPGO. This process containing optimization and recall is repeated until no new loops are recalled.
@@ VII-A Experimental Setup
We first evaluate our proposed spatial bundle adjustment by comparing it with the conventional LiDAR-based bundle adjustment methods BALM2 [24] and HBA [25] (Section VII-B). Next, we assess the localization performance of our complete multi-session merging framework against DCL-SLAM [27] and LAMM [29] (Section VII-C). Moreover, the mapping quality evaluation is detailed in VII-D. An ablation study is then conducted to analyze the contributions of individual components within our framework (Section VII-E). Finally, we test the scalability of our method in handling a large number of sessions (Section VII-F).
@@ VII-B Single-Robot Study
VII-C Multi-Robot Localization Study
@@
VII-D Multi-Robot Mapping Quality Evaluation
In this section, we evaluate the mapping quality of multi-robot using the MapEval [52]. For the MARS-LVIG dataset, the ground-truth map is a high-precision point cloud generated by the DJI L1 LiDAR sensor and processed with the DJI Terra system. For the S3E dataset, since no ground-truth map is available and the ground-truth poses do not contain rotational information, we evaluate the geometric quality of the reconstructed maps using the average plane thickness and planarity metrics. Detailed definitions of the evaluation metrics are provided in [52].
Table V presents the quantitative comparison of mapping quality between our method and LAMM. Our approach achieves the best performance across all datasets, including lower average Wasserstein distance (AWD), lower Chamfer distance (CD), smaller Spatial Consistency Score (SCS), and lower Mean Map Entropy (MME). These results indicate that our method produces higher-quality maps with better alignment to the ground-truth maps. In these three large-scale environments covering areas of over 100,000 square meters, our AWD ranges from 0.25 m to 0.65 m, and the CD ranges from 0.43 m to 1.11 m, demonstrating strong consistency with the ground-truth maps. Furthermore, the lower SCS and MME values suggest that our method achieves better spatial consistency than LAMM, which can be attributed to the proposed spatial BA and the last PGO for improving local accuracy and global consistency. Furthermore, Fig 12 further visualizes the Wasserstein distance distribution on the Island dataset. As shown in (a), the errors are concentrated within the bound (0.93 m), with most voxels around 0.25 m. (b) shows that almost all voxels have Wasserstein distance values below 0.5 m and exhibit a spatially consistent distribution, indicating that the reconstructed multi-robot map is well aligned with the ground-truth map and maintains high spatial consistency.
Table VI reports the averaged plane thickness and planarity evaluation results over all sequences in the S3E dataset. Our method achieves the best performance compared to DCL-SLAM and LAMM, indicating that the proposed spatial bundle adjustment significantly improves the geometric accuracy of multi-robot mapping and effectively mitigates the multi-robot map divergence commonly observed in PGO-based methods.
| Data | Method | Metrics | |||
| AWD(m) | CD(m) | SCS | MME | ||
| Island | LAMM | 1.30 | 6.44 | 0.50 | -5.51 |
| Ours | 0.30 | 0.43 | 0.49 | -6.13 | |
| Town | LAMM | 1.77 | 10.64 | 0.57 | -6.37 |
| Ours | 0.65 | 1.11 | 0.47 | -6.52 | |
| Airport | LAMM | 0.75 | 36.05 | 0.62 | -6.17 |
| Ours | 0.25 | 0.47 | 0.57 | -6.28 | |
| Metric | DCL-SLAM | LAMM | Ours |
| Thickness(m) | 0.12 | 0.14 | 0.08 |
| Planarity | 0.61 | 0.59 | 0.86 |
VII-E Ablation Study
@@
Table VII shows the number of loops in different loop processing stages, as well as the results of the first pose graph optimization with and without loop recall. The experimental results show that the first PGO achieves smaller errors and better performances when mistakenly rejected loops are recalled. Beyond its impact on PGO, loop recall also introduces additional constraints that benefit subsequent bundle adjustment.
@@
Table VIII summarizes the results, with the best values highlighted in bold. The complete framework consistently achieves the lowest RMSE, validating the effectiveness of combining spatial BA and two PGO steps. For the variant of FPGO + BA, although local BA improves relative accuracy, it may disrupt the continuity of odometry since it only refines loop regions. Consequently, its standalone use can increase RMSE compared to FPGO alone. However, the last PGO integrates both local BA constraints and odometry continuity, propagating local accuracy to the global map and significantly reducing RMSE.
@@
In particular, we visualize the comparison between coarse GICP registration and the proposed fine spatial BA. Fig. 15 (a) shows the result of GICP-based coarse alignment, where only basic submap alignment is achieved, and noticeable misalignments appear in tree trunks, traffic signs, and road boundaries. Fig. 15 (b) shows the result after applying the proposed spatial BA, where submaps become fully consistent with clear edge structures and significantly improved geometric quality. This confirms that spatial BA is a critical component for high-precision multi-robot mapping. While GICP only ensures coarse submap-level alignment, the proposed spatial BA performs fine-grained joint optimization at the pose level, enabling globally consistent and locally accurate map reconstruction. As a result, the originally divergent multi-robot point clouds are refined to achieve highly consistent and precise reconstruction, approaching the geometric fidelity typically observed in short-range single-robot mapping.
VII-F Scalability Study
@@
Table IX summarizes the corresponding results of three cases. Our framework achieves a success rate 100% in all experiments, including the five-session, ten-session, and twenty-session scenarios. These results demonstrate the scalability and robustness of the LEMON-Mapping system when dealing with large numbers of multi-robot map merging.
@@ VIII Conclusion
Appendix: PROOF OF LEMMA
VIII-A Proof of Lemma 1
Proof.
@@ VIII-A Proof of Lemma 1
A standard result on the continuous dependence of polynomial roots on coefficients now implies that each ordered root varies continuously with the coefficients , and hence with itself. Therefore, every eigenvalue function is continuous. ∎
VIII-B Proof of Lemma 2
Proof.
The complexity of joint BA over all poses is
where the three terms correspond to evaluating feature residuals, forming the Hessian, and solving the LM linear system.
Expanding the squared and cubic terms via group sizes yields
Thus joint BA includes not only the per-group contributions and , but also all cross-group interaction terms such as , , and , which dominate when multiple groups have comparable size.
In DBA, only the poses in the current diffusion group remain active while all inner groups are frozen. The BA performed in diffusion step therefore has complexity
Summing across all diffusion steps gives the total DBA cost
which eliminates all cross-group terms present in . Since
DBA is never more expensive than joint BA.
Moreover, when groups are approximately balanced with , we obtain
showing that the dominating quadratic and cubic terms are reduced by factors of order and , respectively. This proves that DBA yields a strictly lower computational cost and achieves substantial practical speedups. ∎
VIII-C Proof of Lemma 3
Proof.
We proceed in two steps.
VIII-C1 Block-extraction lower bound for the measurement noise term.
For each measurement cluster the per-measurement contribution is
which is positive semidefinite because it is of the form . Subtracting the block-diagonal matrix that keeps only the bottom-right block yields
The right-hand side can be written as
hence
Summing over all measurements preserves the PSD ordering, and extracting the bottom-right block gives the key inequality
VIII-C2 Schur-complement ordering for Hessian inverses.
Write the Schur complement of in :
Because , we have
For two strict positive definite matrices it follows that . Applying this to yields
VIII-C3 Combine VIII-C1 and VIII-C2.
The bottom-right block of the joint covariance can be written using the Schur complement inverse :
Using (15) and the inverse-ordering we obtain
where the second PSD inequality follows from left- and right-multiplying the PSD matrix by and noting .
Thus , which proves the stated ordering. ∎